OpenStudy (anonymous):

If a studio chain has 8 region managers who each oversee 8 district managers who each in turn oversee 15 studios. How many total studios are there? 640 720 960 1020 1280

OpenStudy (anonymous):

8*8*15= 64*15= 960

OpenStudy (anonymous):

\[8\times8\times15 = 960\]